OpenStudy (anonymous):

How do you find the sin of 105 degrees using the half-angle formula?

OpenStudy (anonymous):

is it the cosine of 52.5

OpenStudy (anonymous):

\[\sin \theta \div2= \pm \sqrt{1-\cos \theta \div2}\]